OpenStudy (anonymous):

(vii) The cube of a single digit number may be a single digit number.true or false

OpenStudy (anonymous):

true

OpenStudy (anonymous):

1^3=1

OpenStudy (anonymous):

medal plz

OpenStudy (anonymous):

True. Eg, 1^3=1,2^3=8